An object slides down at the top of a hill starting from rest it takes 8 seconds to slide down 100 meters at an angle of 60. Then what will be the coefficient of kinetic friction?

1
Expert's answer
2022-02-20T08:20:06-0500

"a=g(\sin\alpha-\mu\cos\alpha),"

"h=\frac{at^2}2,"

"\mu=\tan\alpha-\frac{2h}{gt^2\cos\alpha}=1.09."